OpenStudy (anonymous):

Solve x2 + 2x + 9 = 0

OpenStudy (rational):

you may use quadratic formula 

OpenStudy (shamim):

If u compare both equations u will get a=1 b=2 c=9

OpenStudy (shamim):

Tel us x=?

OpenStudy (anonymous):

\[\frac{ -2 \pm \sqrt{2^{2}-4(1)(9)} }{ 2(1) } = \frac{ -2\pm \sqrt{-32} }{ 2 }\]

OpenStudy (anonymous):

now I'm not sure about this part, so if i'm wrong, correct me anyone. i just wanna help. \[\frac{ -2\pm \sqrt{-1}\sqrt{32} }{ 2 } = \frac{ -2 \pm (i) (4\sqrt{2}) }{ 2 } = \frac{ -2 \pm 4i \sqrt{2} }{ 2 } = \frac{ 2 (-1 \pm 2i \sqrt{2}) }{ 2 } = -1 \pm 2i \sqrt{2}\]